OpenStudy (anonymous):

2x^2 + x +4 =0 wt=rite quation in standard form then identify a,b, and c

OpenStudy (anonymous):

2x^2+x+4=0 =>x^2+x/2+2=0 and so u get a=1 ,b=1/2 and c=2.

OpenStudy (anonymous):

\[2x ^{2}+x+4=0\] By simplifying this you will \[x ^{2}+x/2+2=0\] You can say it as \[ax ^{2}+bx+c=0\] so a=1, b=1/2 and Constant C=2

OpenStudy (anonymous):

ok how bout a negative like -3x^2 - 2 = x^2 + 3x

OpenStudy (anonymous):

Again you have to simplify it do it as the same practice..if there is any doubts can ping me